# Efficient 6-dimensional phase space reconstruction from experimental measurements using generative machine learning

### v1

@@ Abstract
Next-generation accelerator concepts which hinge on the precise shaping of beam distributions, demand equally precise diagnostic methods capable of reconstructing beam distributions within 6-dimensional position-momentum spaces. However, the characterization of intricate features within 6-dimensional beam distributions using conventional diagnostic techniques necessitates hundreds of measurements, using many hours of valuable beam time. Novel phase space reconstruction techniques are needed to substantially reduce the number of measurements required to reconstruct detailed, high-dimensional beam features in order to resolve complex beam phenomena, and as feedback in precision beam shaping applications. In this study, we present a novel approach to reconstructing detailed 6-dimensional phase space distributions from experimental measurements using generative machine learning and differentiable beam dynamics simulations. We demonstrate that for a collection of synthetic beam distribution test cases that this approach can be used to resolve 6-dimensional phase space distributions using basic beam manipulations and as few as 20 2-dimensional measurements of the beam profile, without the need for prior data collection or model training. We also demonstrate an application of the reconstruction method in an experimental setting at the Argonne Wakefield Accelerator, where it is able to reconstruct the beam distribution and accurately predict previously unseen measurements 75x faster than previous methods.
@@ I Introduction
Current and future particle accelerators rely on the precise control of beam distributions in six-dimensional position-momentum phase space, especially as accelerators push the boundaries of intensity and brightness for collider and light source applications [1]. For example, magnetized electron beams for hadron beam cooling [2], two bunch operations in free electron lasers [3], and drive/witness bunch pairs for high efficiency wakefield acceleration [4], are enabled by the precise control of 6-dimensional phase spaces beyond macroscopic (RMS) beam properties.
@@
Achieving this level of control requires measurement techniques that provide detailed information about the beam distribution in all 6 phase space coordinates, including cross-correlations between the different phase space coordinates. Measurements of this type are needed to provide feedback to accelerator operators or autonomous control algorithms to tune accelerator parameters such that the measured beam distribution matches what is needed by applications. Additionally, detailed characterization of 6-dimensional phase space distributions is necessary to resolve complex beam dynamics phenomena that couples particle motion along multiple axes, such as coherent synchrotron radiation [5] or plasma wakefield accelerators [6].
@@
A wide variety of beam manipulation and diagnostic techniques have been developed to measure and predict detailed characteristics of phase space distributions. These techniques can involve rotating the beam in phase space (tomography) [7, 8, 9, 10], using masks or meshes to isolate and observe the dynamics of portions of the transverse beam [11], or using specialized, non-destructive beam diagnostics such as laser wires [12], or machine learning models [13, 14, 15] to predict the beam distribution from experimental measurements.
@@
Reconstructing detailed 5- or 6-dimensional phase space distributions from experimental data has proven to be substantially more difficult than reconstructing lower dimensional spaces (4 or less phase space coordinates). Experimental 5-dimensional phase space characterization has been done in limited instances, once at the ARES beamline with a polarizable X-band transverse deflecting cavity [16, 17] using the Simultaneous Algebraic Reconstruction Technique (SART) algorithm [18], and once at the Spallation Neutron Source (SNS) using a set of movable masking slits [19]. Full 6-dimensional reconstruction of a single beam distribution has also only been done once, by combining multiple, scanning masking slits with a dipole spectrometer and a bunch shape monitor at the SNS beamline [20]. However, these measurements required a significant amount beam time resources to carry out (960 measurements over $\sim 28$ hrs for the 5-dimensional ARES case, over 5 million measurements over $\sim 32$ hrs for the 6-dimensional SNS measurement), making these measurement procedures impractical for regular use as feedback for online accelerator tuning or for understanding complex beam dynamics phenomena. Reconstruction methods that significantly reduce the number of measurements needed to reconstruct the 6-dimensional beam distribution in order to be used during regular accelerator operations.
@@
In previous work, we introduced and demonstrated a novel method for reconstructing detailed beam distributions using generative machine learning models and differentiable beam dynamics simulations [21], which we refer to here as generative phase space reconstruction (GPSR). The generative machine learning model creates nearly arbitrary beam distributions in 6-dimensional phase space by transforming random samples drawn from a fixed probability distribution into macroparticles with realistic position-momentum coordinates using a small neural network. A differentiable beam dynamics simulation, which allows for analytical calculation of derivatives during each computational step, is then used to train parameters of the generative model (which in turn modifies the generated beam distribution) by minimizing the difference between experimental measurements and simulated predictions. Combining generative models with differentiable beam dynamics simulations allows the generative model to be trained solely from experimental measurements in a short amount of time, removing the need for large initial training sets usually required by machine learning workflows. This method was used to reconstruct detailed, correlated, 4-dimensional phase space distributions using a single quadrupole scan (10 measurements of the beam distribution) by taking advantage of the information contained in conventional measurements of the transverse beam profile using a YAG screen 111noting that a quadrupole scan does not provide information about the longitudinal phase space distribution. More recently, this method was also used to reconstruct beams with large angular momentum components (magnetization) and large emittance ratios [23].
@@
In this work, we extend the use of GPSR to reconstructing 6-dimensional phase space distributions from experimental measurements. We demonstrate in simulation that a diagnostic beam line with a single quadrupole, transverse deflecting cavity, and dipole spectrometer can be used to resolve detailed characteristics of 6-dimensional phase space distributions using as little as 20 measurements, reducing the beam time necessary to reconstruct the beam distribution from many hours to several minutes. We show that the reconstruction technique accurately reconstructs a variety of different beam distributions including pure Gaussian distributions, correlated distributions, and beam distributions similar to those produced by the emittance exchange (EEX) beamline at the Argonne Wakefield Accelerator [24]. We then apply the GPSR algorithm to reconstructing a beam distribution from experimental measurements and demonstrate that is accurately predicts additional measurements in experiment. Finally, we discuss the current limitations and advantages of the reconstruction technique.
@@ II Methods
The method we use to reconstruct high-dimensional phase space distributions targets two issues encountered by conventional reconstruction techniques.
@@
First, these measurement techniques use low-dimensional projections of the 6-dimensional beam to infer distribution structure. Conventional algebraic reconstruction techniques such as Maximum Entropy Tomography (MENT) [25] and SART typically use 1-dimensional projections of screen images to produce a higher dimensional reconstruction. Alternatively, measuring the intensity of the beam distribution through a set of slits, as done in [20], further reduces the dimensionality of the measurement down to a scalar quantity. As a result, these methods lose most of the information about the beam distribution by projecting them to lower-dimensional measurements, requiring significantly more individual measurements of the beam distribution to resolve higher-dimensional features.
@@
Second, conventional techniques for describing beam distributions scale extremely poorly when describing 6-dimensional beam distributions. Both of the methods listed above use a high-dimensional histogramming approach to describe the beam distribution, i.e., they solve for the beam density inside a number of high-dimensional voxels along a 6-dimensional mesh. While this formalism is has been successfully used to describe 1-2 dimensional distributions, it becomes extremely expensive to describe beam distributions in 6-dimensional space as the number of bins grows exponentially with the number of dimensions. For example, resolving a beam distribution with 100 bins per dimension results in $10^{12}$ voxels needed to describe the full 6-dimensional distribution. Alternative methods for representing beam distributions, such as distribution moments or principal component analysis techniques, suffer from similar dimensionality issues.
To overcome these challenges we use two recent developments in machine learning and differentiable beam dynamics modeling.
@@
First, we use a generative machine learning model to represent a distribution of macro particles in 6-dimensional space, inspired by the development of a technique known as neural radiance fields [26], which uses neural networks to represent mass and color density functions in 3-dimensional space. This model uses a small neural network to transform randomly generated samples from a 6-dimensional multivariate Normal distribution into real macroparticle phase space coordinates in position-momentum space. Neural networks of sufficient complexity are universal function approximators [27], enabling this transformation network to produce particle distributions with nearly arbitrary structure in high-dimensional phase space. For the work here, we found that a fully connected neural network with 2 layers, 50 neurons each, connected by Tanh activation functions was sufficient to represent nearly arbitrary beams with high enough detail. With this method, the distribution of macroparticles in 6-dimensional phase space is entirely controlled by the parameters of the neural network, resulting in a parameterization using approximately 1000 free parameters, as opposed to the millions of parameters needed for mesh-based representations of 6-dimensional distributions.
@@
The reconstruction process determines the parameters of the neural network at the center of the generative beam model by solving an optimization problem, shown in Fig. 1. Optimization is used to minimize a loss function that is proportional to the mean squared discrepancy between simulated measurement predictions and experimental measurements, in this case the per-pixel intensity of a transverse YAG screen. Due to the number of free parameters contained inside the generative beam distribution model, solving this optimization problem in a reasonable amount of time is beyond the capabilities of black box optimization algorithms. Black box optimization algorithms, such as Bayesian optimization [28], use evaluations of the loss function to iteratively solve for the ideal parameter set that reaches the minimum objective value. However, due to the so-called “curse-of-dimensionality” [29], these algorithms require an exponentially increasing number of iterations to reach a minimum value when so many free parameters are present.
@@
Given a large number of free parameters, it is necessary to use gradients of the loss function in addition to the loss function value to inform optimization and use gradient-descent-based algorithms to solve optimization problems. Unfortunately, this is also prohibitively difficult when using conventional beam dynamics simulations to predict experimental measurements, since calculating the gradients numerically in these cases requires finite-difference methods that also scale poorly with the number of free parameters used in optimization.
@@
To address this challenge, we take advantage of recent developments in automatic differentiation programming to substantially reduce the cost of evaluating gradients of simulation outputs with respect to many input parameters. Automatic differentiation is the technique of tracking derivative information alongside each computation step during physics simulations. This in turn, allows analytical evaluation of simulation output derivatives by using the chain rule in reverse to determine derivatives with respect to input parameters, as process commonly referred to as “backpropagation” [30] or “adjoint-differentiation” [31]. Computational costs of calculating the derivatives in this manner is roughly equivalent to the cost of evaluating the simulation itself, and more importantly, is independent of the number of input parameters the derivative is taken with respect to, thus making it substantially cheaper to compute than finite-difference methods when optimizing with respect to a large number of input parameters.
It should be noted that this process is similar to, but distinct from, previous uses of automatic differentiation in accelerator physics, often referred to as “forward-mode” differentiation [31], “differential algebra” [32], or, when computing higher derivatives ,“truncated power series algebra” [33]. These techniques are well suited for calculating particle transport dynamics up to arbitrary order, however they do not scale well to calculating gradients with respect to thousands of input parameters.
@@
Facilitating the use of backpropagation in the context of phase space reconstruction requires beam dynamics and diagnostic simulations that support the tracking of derivatives during evaluations. To this end, we have developed the simulation package Bmad-X [34] which re-implements beam transport through a number of simple accelerator elements as simulated by the software package Bmad [35]. Bmad-X implements the base mathematics of beam transport in Python using a package-agnostic approach, enabling arbitrary numerical computation packages to be used to do the underlying computation. This work uses the machine learning library PyTorch [36], which implements backpropagation, to perform beam dynamics computations. Additionally, we simulate the measurement of transverse beam profile intensity using a screen diagnostic by using kernel density estimation (KDE) [37] as opposed to normal histogramming to preserve differentiability. By using differentiable beam dynamics and diagnostic simulations we are able to cheaply compute derivatives for use in gradient descent optimization of beam distribution parameters to reconstruct the beam distribution. More recently, other, similar beam dynamics simulation packages have been developed, namely the Cheetah python package [38], although they were not used in this work.
@@
In previous work [21], we demonstrated that this reconstruction technique is able to produce accurate predictions of the beam distribution using images gathered from a single quadrupole scan. To extend this work towards resolving 6-dimensional phase spaces, we add a transverse deflecting cavity (TDC) and a dipole spectrometer to the diagnostic beamline, shown in Fig. 1. In this case, quadrupole Q4 is scanned to measure the transverse phase space distribution, where quadrupoles Q1-Q3 are used to focus the beam onto YAG1 before Q4 is scanned. Focusing the beam on the diagnostic screen improves the measurement resolution and increases the range of Q4 strengths that can be scanned over, since the transverse beam profile needs to be kept within a region of interest on both of the diagnostic screens. Pairing a TDC that kicks the beam vertically to resolve the current profile of the beam, with a horizontally bending dipole magnet which measures beam energy spread, is a common approach taken to measure the longitudinal phase space distribution [39]. The diagnostic setup used in this work is motivated by the notion that combining transverse information from the quadrupole scan with the longitudinal phase space manipulations should provide enough information to resolve the full 6-dimensional phase space distribution.
We demonstrate the 6-dimensional reconstruction technique using a simulation of the 6-dimensional reconstruction diagnostic beamline with similar parameters to the diagnostic beamline used at the Argonne Wakefield Accelerator (AWA). The beam dynamics and screen diagnostics were simulated in the differentiable beam dynamics package Bmad-X. Transverse diagnostic screens YAG1 and YAG2 are placed along each beam path when the dipole is off and on respectively. These simulated screens have a region of interest that is 5 x 5 mm (or 200 x 200 pixels) in size, with a resolution of 25 $\mu$m/px. The L-band transverse deflecting cavity used in simulation (1.3 GHz, $L=0.48$ m) operated with a peak field in a range from 0-3 MV depending on the beam distribution, consistent with the operational range of the AWA deflecting cavity [40]. The rectangular dipole spectrometer ($L=0.3018$ m) has a bend angle of 20 degrees.

### v2

# Efficient 6-dimensional phase space reconstructions from experimental measurements using generative machine learning

@@ Abstract
Next-generation accelerator concepts, which hinge on the precise shaping of beam distributions, demand equally precise diagnostic methods capable of reconstructing beam distributions within 6-dimensional position-momentum spaces. However, the characterization of intricate features within 6-dimensional beam distributions using current diagnostic techniques necessitates a substantial number of measurements, using many hours of valuable beam time. Novel phase space reconstruction techniques are needed to reduce the number of measurements required to reconstruct detailed, high-dimensional beam features in order to resolve complex beam phenomena, and as feedback in precision beam shaping applications. In this study, we present a novel approach to reconstructing detailed 6-dimensional phase space distributions from experimental measurements using generative machine learning and differentiable beam dynamics simulations. We demonstrate that this approach can be used to resolve 6-dimensional phase space distributions from scratch, using basic beam manipulations and as few as 20 2-dimensional measurements of the beam profile. We also demonstrate an application of the reconstruction method in an experimental setting at the Argonne Wakefield Accelerator, where it is able to reconstruct the beam distribution and accurately predict previously unseen measurements 75x faster than previous methods.
@@ I Introduction
Current and future particle accelerators rely on the precise control of beam distributions in 6-dimensional position-momentum phase space, especially as accelerators push the boundaries of intensity and brightness for collider and light source applications [1]. For example, magnetized electron beams for hadron beam cooling [2], two bunch operations in free electron lasers [3], and drive/witness bunch pairs for high efficiency wakefield acceleration [4], are enabled by the precise control of 6-dimensional phase spaces beyond macroscopic (RMS) beam properties.
@@
Achieving this level of control requires measurement techniques that provide detailed information about the beam distribution in all 6 phase space coordinates, including cross-correlations between the different phase spaces. Measurements of this type are needed to provide feedback to accelerator operators or autonomous control algorithms to tune accelerator parameters such that the measured beam distribution matches what is needed by applications. Additionally, detailed characterization of 6-dimensional phase space distributions is necessary to resolve complex beam dynamics phenomena that couples particle motion along multiple axes, such as coherent synchrotron radiation [5] or plasma wakefield accelerators [6].
@@
A wide variety of beam manipulation and diagnostic techniques have been developed to measure and predict detailed characteristics of phase space distributions. These techniques can involve rotating the beam in phase space (tomography) [7, 8, 9, 10], using masks or meshes to isolate and observe the dynamics of portions of the transverse beam [11], using specialized, non-destructive beam diagnostics such as laser wires [12], or using machine learning models [13, 14, 15] to predict the beam distribution from experimental measurements.
@@
Reconstructing detailed 5- or 6-dimensional phase space distributions from experimental data has proven to be substantially more difficult than reconstructing lower dimensional spaces (4 or less phase space coordinates). Experimental 5-dimensional phase space characterization has been done in limited instances, once at the ARES beamline with a polarizable X-band transverse deflecting cavity [16, 17] using the Simultaneous Algebraic Reconstruction Technique (SART) algorithm [18], and once at the Spallation Neutron Source (SNS) using a set of movable masking slits [19]. Full 6-dimensional reconstruction of a single beam distribution has also only been done once, by combining multiple, scanning masking slits with a dipole spectrometer and a bunch shape monitor at the SNS beamline [20]. However, these measurements required a significant amount beam time resources to carry out (960 measurements over $\sim 28$ hrs for the 5-dimensional ARES case, $\sim 5$ million measurements over $\sim 32$ hrs for the 6-dimensional SNS measurement), making these measurement procedures impractical for regular use as feedback for online accelerator tuning or for understanding complex beam dynamics phenomena. There is a need for reconstruction methods that significantly reduce the number of measurements required to reconstruct the 6-dimensional beam distribution in order to be used during regular accelerator operations.
@@
In previous work, we introduced and demonstrated a novel method for reconstructing detailed beam distributions, which we refer to here as generative phase space reconstruction (GPSR), using generative machine learning models and differentiable beam dynamics simulations [21]. The generative machine learning model creates nearly arbitrary beam distributions in 6-dimensional phase space by transforming random samples drawn from a fixed probability distribution into macroparticles with realistic position-momentum coordinates. A differentiable beam dynamics simulation, which allows for analytical calculation of derivatives during each computational step [22], is then used to train parameters of the generative model (which in turn modifies the generated beam distribution) by minimizing the difference between experimental measurements and simulated predictions. Combining generative models with differentiable beam dynamics simulations allows the generative model to be trained solely from experimental measurements in a short amount of time, removing the need for large initial training sets usually required by machine learning workflows. This method was used to reconstruct detailed, correlated, 4-dimensional phase space distributions using a single quadrupole scan (10 measurements of the beam distribution) by taking advantage of the information contained in conventional measurements of the transverse beam profile using a YAG screen 111noting that a quadrupole scan does not provide information about the longitudinal phase space distribution. More recently, this method was also used to reconstruct beams with large angular momentum components (magnetization) and large emittance ratios [24].
@@
In this work, we extend the use of GPSR to reconstructing 6-dimensional phase space distributions from experimental measurements. We demonstrate in simulation that a diagnostic beamline consisting of quadrupoles, a transverse deflecting cavity, and a dipole spectrometer can be used to resolve detailed characteristics of 6-dimensional phase space distributions using as few as 20 measurements. This reduces the time necessary to reconstruct the beam distribution from many hours to several minutes. We show that the reconstruction technique accurately reconstructs a variety of different beam distributions including correlated Gaussian distributions, nonlinear distributions, and beam distributions similar to those produced by the emittance exchange (EEX) beamline at the Argonne Wakefield Accelerator (AWA) [25]. We then apply the GPSR algorithm to reconstructing a beam distribution from experimental measurements at AWA and demonstrate that it makes accurate predictions of previously unseen measurements. Finally, we discuss current limitations and advantages of the reconstruction technique.
@@ II Methods
The method we use to reconstruct 6-dimensional phase space distributions addresses two issues encountered by conventional reconstruction techniques.
@@
First, conventional reconstruction techniques often use 1 or 0-dimensional (scalar) projections of the 6-dimensional phase space to infer features of the distribution. Conventional algebraic reconstruction techniques such as Maximum Entropy Tomography (MENT) [26] and SART [18] typically use 1-dimensional projections of screen images to produce a higher dimensional reconstruction. Measuring the intensity of the beam distribution through a set of slits, as done in [20], further reduces the measurement down to a scalar quantity. As a result, these methods lose a significant amount of information about the beam distribution, requiring more individual measurements of the distribution to resolve 6-dimensional features. The GPSR technique enables us to fully utilize the detailed information contained in 2-dimensional images of the beam distribution, without the need to project to lower dimensions.
@@
Second, conventional techniques for representing the distribution of particles in a beam scale poorly when describing 6-dimensional distributions. Both of the methods listed above use a high-dimensional histogramming approach to describe the beam distribution, i.e., they solve for the beam density inside a number of high-dimensional voxels along an N-dimensional mesh. While this formalism is used to effectively describe distributions in 1-2 dimensions, it becomes impractical to describe beam distributions in 6-dimensional space this way, as the number of bins grows exponentially with the number of dimensions. For example, resolving a beam distribution with 100 bins per dimension results in $10^{12}$ voxels needed to describe the full 6-dimensional distribution. Alternative methods for representing beam distributions, such as distribution moments or principal component analysis techniques, reduce reconstruction detail.
To address these challenges, GPSR introduces two new concepts shown in Fig. 1 to describe and solve for detailed 6-dimensional phase space distributions.
@@
First, we use a generative machine learning model to represent a distribution of macro particles in 6-dimensional phase space. This is inspired by the development of neural radiance fields [27], which uses neural networks to represent mass and color density functions in 3-dimensional space. GPSR uses a generative model, in this case a small neural network, to transform randomly generated samples from a 6-dimensional multivariate Normal distribution into real macroparticle phase space coordinates. Neural networks of sufficient complexity are universal function approximators [28], enabling the generative model to produce particle distributions with nearly arbitrary structure in 6-dimensional phase space. For the work here, we found that a fully connected neural network with 2 layers, 20 neurons each, connected by Tanh activation functions, was sufficient to represent nearly arbitrary beams with high enough detail. With this method, the distribution of macroparticles in phase space is entirely controlled by the parameters of the neural network, resulting in a parameterization of the distribution using approximately 1000 free parameters, as opposed to the millions of parameters needed for mesh-based representations of 6-dimensional distributions.
@@
The reconstruction process determines the parameters of the generative beam model by solving an optimization problem, where the goal is to minimize the mean squared error between simulated predictions and experimental measurements, in this case the per-pixel intensity of the transverse YAG screens. Due to the number of free parameters contained inside the generative beam distribution model, solving this optimization problem in a reasonable amount of time is beyond the capabilities of black box optimization algorithms because of the so-called “curse of dimensionality” [29].
@@
Given a large number of free parameters, it is necessary to use gradient-descent-based algorithms to solve the GPSR optimization problem. Unfortunately, this is also prohibitively difficult when using conventional beam dynamics simulations to predict experimental measurements, since calculating the gradients numerically in these cases requires finite-difference methods that also scale poorly with the number of free parameters used in optimization.
@@
To address this challenge, GPSR leverages “backwards-mode” [30] automatic differentiation to substantially reduce the cost of evaluating gradients of simulation outputs with respect to input parameters. Automatic differentiation is the technique of tracking derivative information alongside each computation step during physics simulations. This in turn, allows analytical evaluation of simulation output derivatives by using the chain rule in reverse to determine derivatives, as process commonly referred to as “backpropagation” [30] or “adjoint-differentiation” [31]. Computational costs of calculating the derivatives in this manner is roughly equivalent to the cost of evaluating the simulation itself, and more importantly, is independent of the number of input parameters the derivative is taken with respect to. This makes calculating derivatives substantially cheaper to compute than finite-difference methods when optimizing with respect to a large number of input parameters. It should be noted that this process is similar to, but distinct from, previous uses of automatic differentiation in accelerator physics, often referred to as “forward-mode” differentiation [31], “differential algebra” [32], or, when computing higher derivatives ,“truncated power series algebra” [33]. These techniques are well suited for calculating particle transport dynamics up to arbitrary order, however they do not scale well to calculating gradients with respect to thousands of input parameters.
@@
Facilitating the use of backpropagation in the context of GPSR requires beam dynamics and diagnostic simulations that support the tracking of derivatives during evaluations. To this end, we have developed the simulation package Bmad-X [22] which re-implements beam transport through a number of simple accelerator elements using the machine learning library PyTorch [35] which implements backpropagation. Additionally, we simulate the measurement of transverse beam profile intensity on a screen diagnostic by using kernel density estimation (KDE) [36] as opposed to normal histogramming to preserve differentiability. By using differentiable beam dynamics simulations, we are able to cheaply compute derivatives for use in gradient descent optimization of beam distribution parameters to reconstruct the beam distribution. More recently, similar beam dynamics simulation packages have been developed, namely the Cheetah Python package [37], although they were not used in this work.
@@
In previous work [21], we demonstrated that this reconstruction technique is able to produce accurate predictions of the 4-dimensional beam distribution using images gathered from a single quadrupole scan. To extend this work towards resolving 6-dimensional phase spaces, we add a transverse deflecting cavity (TDC) and a dipole spectrometer to the diagnostic beamline shown in Fig. 1. In this case, quadrupoles Q1-Q3 are used to focus the beam onto YAG1 and quadrupole Q4 is scanned to measure the transverse phase space distribution. Focusing the beam on the diagnostic screen improves the measurement resolution and increases the range of Q4 strengths that can be scanned over since the transverse beam profile needs to be kept within a region of interest on both of the diagnostic screens. Pairing a TDC that kicks the beam vertically to resolve the current profile of the beam with a horizontally bending dipole magnet, which measures beam energy spread, is a common approach taken to measure the longitudinal phase space distribution [38]. The diagnostic setup used in this work is motivated by the notion that combining transverse information from the quadrupole scan with the longitudinal phase space manipulations should provide enough information to resolve the full 6-dimensional phase space distribution.
We demonstrate the 6-dimensional reconstruction technique using a simulation of the diagnostic beamline at AWA. Transverse diagnostic screens YAG1 and YAG2 are placed along each beam path when the dipole is off and on respectively. These simulated screens have a region of interest that is 5 x 5 mm (or 200 x 200 pixels) in size, with a resolution of 25 $\mu$m/px. The L-band transverse deflecting cavity used in simulation (1.3 GHz, $L=0.48$ m) operated with a peak field in a range from 0-3 MV depending on the beam distribution, consistent with the operational range of the AWA deflecting cavity [39]. The rectangular dipole spectrometer ($L=0.3018$ m) has a bend angle of 20 degrees.